# Overparameterized Multiple Linear Regression as Hyper-Curve FittingE. Atza and N. Budko

### v1

@@ 5 Removal of improper predictors
From the analysis and examples of the previous sections it is clear that a linear model, be it the classical MLR implementation or the present IR formulation, will produce reasonably exact predictions even if some of the predictor variables do not satisfy the model assumptions. In the regularized IR formulation of Section 4 the model assumption is that the predictor is a polynomial in the dependent variable $y$ of degree at most $r^{*}-1$, where $r^{*}$ is the rank of the optimally regularized $\hat{A}$. In the case of exact data discussed in Section 3, the training set may become complete way before some of the predictor functions have been properly sampled. Success of predictions in the presence of such ‘improper’ predictors gives an illusion of understanding of the underlying natural phenomena [21]. In Section 4 we have also investigated the case where the noise is present only in some of the predictors. Such noisy predictors may be negatively affecting the predictive power of the model. Thus, there appear to be two good reasons to detect and possibly discard both the high-degree/non-functional and the noisy predictors from the model. Additionally, in biological and agricultural applications, the large number of predictors included in the FOP dataset is often due to a broad (untargeted) experimental search in the absence of any a-priori information. The practical goal of such studies is to identify a preferably small subset of microbiota, fungi, molecules, metabolites, or genes, allowing for future targeted and less expensive measurements of these predictor variables.
@@
Fig. 4 presents two examples of the feature removal error on the training ${\bm{y}}$-data (solid blue) together with the corresponding non-accessible feature removal errors on the test ${\bm{y}}_{\rm t}$-data (dashed gray). In the left plot, where the level of noise is $\sigma=0.05$, the choice of the optimal threshold $\tau_{\rm opt}$ from the minimum of the training ${\bm{y}}$-data corresponds to the smallest possible number of predictors. In the right plot, where $\sigma=0.1$, the detected minimum of the feature removal error is situated in the second valley of the error function and corresponds to many unnecessary predictors retained in the model, as the depth of both valleys is approximately the same. Checking for other local minima in this case and choosing the leftmost along the $\tau$-axis would be a better strategy to find $\tau_{\rm opt}$. Regardless, it is comforting to observe that the feature removal errors on the training and test datasets appear synchronous in these numerical examples.
@@
In Fig. 5 (top) the green triangles show the true ‘degrees’ (lengths of the nonzero parts of the corresponding columns of the coefficient matrix $A$, used to generate the predictor data) for each predictor. The vertical dashed dark-grey lines indicate the predictors containing the Gaussian noise. The vertical dashed red lines indicate the non-functional predictors. The horizontal solid orange line depicts the optimal degree $r^{*}$.
In Fig. 5 (bottom) the orange circles show the value of the column prediction error $\rho({\bm{x}}_{j})$ on the test $X_{\rm t}$-data for each predictor obtained with $r=r^{*}$. The dashed blue line is the optimal threshold $\tau_{\rm opt}$. All predictors with the errors above that line (solid light-gray vertical lines) are discarded. Notice that all noisy predictors, as well as both non-functional predictors are successfully removed.
@@
Diamond markers in Fig. 3 show by how much the removal of ‘improper’ predictors improves the prediction error $\rho({\bm{y}}_{t})$. In fact, it is close to the minimal achievable error with all predictors retained. That level of error was, however, not always obtained with the crude global regularization procedure. Hence, the removal of predictors may sometimes compensate for the errors in determining the optimal regularization parameter.
@@ 6 Application to chemometric data
In this section we present the application of the regularized IR model with predictor removal to the experimental dataset that is often used to compare various regularized MLR models, such the Principal Component Regression (PCR) or the Partial Least Squares (PLS) [19]. This dataset was first published in [23] and is also included in the $R$-library pls [22] under the name of yarn. The dataset consists of $28$ Near InfraRed (NIR) spectra of Polyethylene terephthalate (PET) yarns, measured at $p=268$ wavelengths, and $28$ corresponding yarn densities. Seven of these $28$ samples are designated as the test dataset and the remaining $n=21$ constitute the training dataset. The predictor variables (spectral amplitudes) change continuously with the wavelength, which becomes clear if one plots the rows of $X$ as curves, see the middle plot of Fig. 5(a).
@@
The training dataset columns have been sorted by the ${\bm{y}}$-data. Since $n=21<150$, we use the visual inspection of the fitted polynomials, see Fig. 5(b), rather than the CV procedure to determine the optimal degree $r^{*}=5$. The corresponding validation error $\rho(X_{\rm v})$ is shown in Fig. 6(a) (left plot) with its minimum reached at the maximum displayed degree $r^{*}=5$. The first three rows of the coefficient matrix $\tilde{A}$ obtained with $r=r^{*}$ are shown as curves in the bottom plot of Fig. 5(a). These curves correspond to the constant, linear and quadratic components in the polynomial fits $x_{j}(y)$ of each predictor. The column-wise prediction errors $\chi_{j}$ are shown in the top plot of Fig. 5(b) (solid blue), together with the optimal threshold $\tau_{\rm opt}$ (horizontal dashed red). The $\rho({\bm{y}})$ error with $r=r^{*}=5$ as a function of $\tau$ can be seen in Fig. 6(a) (right plot, solid blue), with its minimum indicated by the red square marker. This minimum is close to the minimum of the inaccessible $\rho({\bm{y}}_{\rm t})$ error (dashed gray). Here, as in the numerical examples of Section 5, we also observe the synchronous behavior of the feature removal errors on the training and test datasets.
@@
The removed ‘improper’ predictors are indicated with vertical solid light-gray lines in Fig. 5(b) (middle and bottom). We have also extracted a few typical predictor data (colored vertical lines in Fig. 5(a)) and displayed them in Fig. 5(b) using the same colors and line styles for the corresponding polynomial fits. The retained predictors are in the left plot of Fig. 5(b) and the removed predictors are in the right plot. It is easy to recognize the cone-type data, similar to the data in Fig. 1 (right), in one of the discarded predictors (blue). All predictor data in the discarded gray zone around the dashed blue line in Fig. 5(a) have this ‘conical’ shape. This means that the NIR data in this frequency band is situated on a hyper-cone. Even if this frequency band is not required to build a high-quality predictive IR model, it may become useful for testing other higher-dimensional manifold and nonlinear models.
@@
Finally, in Fig. 6(b) we display the standard quality-of-prediction scatter plot which compares the exact values of ${\bm{y}}$ against their predictions $\hat{\bm{y}}$. As can be seen, the removal of the ‘improper’ predictors significantly improves the quality of prediction. In fact, the prediction error $\rho({\bm{y}}_{\rm t})$ on the test dataset goes down from $0.28$ to $0.05$ after the predictor removal procedure.

### v2

@@ 5 Removal of improper predictors
From the analysis and examples of the previous sections it is clear that a linear model, be it the classical MLR implementation or the present IR formulation, will produce reasonably exact predictions even if some of the predictor variables do not satisfy the model assumptions. In the regularized IR formulation of Section 4 the model assumption is that the predictor is a polynomial in the dependent variable $y$ of degree at most $r^{*}-1$, where $r^{*}$ is the rank of the optimally regularized $\hat{A}$. In the case of exact data discussed in Section 3, the training set may become complete way before some of the predictor functions have been properly sampled. Success of predictions in the presence of such ‘improper’ predictors gives an illusion of understanding of the underlying natural phenomena [12].
In Section 4 we have also investigated the case where the noise is present only in some of the predictors. Such noisy predictors may be negatively affecting the predictive power of the model. Thus, there appear to be two good reasons to detect and possibly discard both the high-degree/non-functional and the noisy predictors from the model. Additionally, in biological and agricultural applications, the large number of predictors included in the FOP dataset is often due to a broad (untargeted) experimental search in the absence of any a-priori information. The practical goal of such studies is to identify a preferably small subset of microbiota, fungi, molecules, metabolites, or genes, allowing for future targeted and less expensive measurements of these predictor variables.
@@
Figure 5 presents two examples of the feature removal error on the training $\bm{y}$}-data (solid blue) together with the corresponding non-accessible feature removal errors on the test ${\bm{y}}_{\rm t}$-data (dashed gray). In the left plot, where the level of noise is $\sigma=0.05$, the choice of the optimal threshold $\tau_{\rm opt}$ from the minimum of the training $\bm{y}$}-data corresponds to the smallest possible number of predictors. In the right plot, where $\sigma=0.1$, the detected minimum of the feature removal error is situated in the second valley of the error function and corresponds to many unnecessary predictors retained in the model, as the depth of both valleys is approximately the same. Checking for other local minima in this case and choosing the leftmost along the $\tau$-axis would be a better strategy to find $\tau_{\rm opt}$. Regardless, it is comforting to observe that the feature removal errors on the training and test datasets appear synchronous in these numerical examples.
@@
In Figure 6 (top) the green triangles show the true ‘degrees’ (lengths of the nonzero parts of the corresponding columns of the coefficient matrix $A$, used to generate the predictor data) for each predictor. The vertical dashed dark-grey lines indicate the predictors containing the Gaussian noise. The vertical dashed red lines indicate the non-functional predictors. The horizontal solid orange line depicts the optimal degree $r^{*}$.
In Figure 6 (bottom) the orange circles show the value of the column prediction error $\rho({\bm{x}}_{j})$ on the test $X_{\rm t}$-data for each predictor obtained with $r=r^{*}$. The dashed blue line is the optimal threshold $\tau_{\rm opt}$. All predictors with the errors above that line (solid light-gray vertical lines) are discarded. Notice that all noisy predictors, as well as both non-functional predictors are successfully removed.
@@
Diamond markers in Figure 3 and Figure 4 show by how much the removal of ‘improper’ predictors improves the prediction error $\rho({\bm{y}}_{t})$. In fact, it is close to the minimal achievable error with all predictors retained. That level of error was, however, not always obtained with the present crude global regularization procedure. Hence, the removal of predictors may sometimes compensate for the errors in determining the optimal regularization parameter.
@@ 6 Application to chemometric data
In this section we present the application of the regularized IR model with predictor removal to the experimental dataset that is often used to compare various regularized MLR models, such as the Principal Component Regression (PCR) or the Partial Least Squares (PLS) [9]. This dataset was first published in [18] and is also included in the $R$-library pls [13] under the name of yarn. The dataset consists of $28$ Near InfraRed (NIR) spectra of Polyethylene terephthalate (PET) yarns, measured at $p=268$ wavelengths, and $28$ corresponding yarn densities. Seven of these $28$ samples are designated as the test dataset and the remaining $n=21$ constitute the training dataset. The predictor variables (spectral amplitudes) change continuously with the wavelength, which becomes clear if one plots the rows of $X$ as curves, see the middle plot of Figure 7(a).
@@
The training dataset columns have been sorted by the $\bm{y}$}-data. Since $n=21<150$, we use the visual inspection of the fitted polynomials, see Figure 7(b), rather than the CV procedure to determine the optimal degree $r^{*}=5$. The corresponding validation error $\rho(X_{\rm v})$ is shown in Figure 8(a) (left plot) with its minimum reached at the maximum displayed degree $r^{*}=5$. The first three rows of the coefficient matrix $\tilde{A}$ obtained with $r=r^{*}$ are shown as curves in the bottom plot of Figure 7(a). These curves correspond to the constant, linear and quadratic components in the polynomial fits $x_{j}(y)$ of each predictor. The column-wise prediction errors $\chi_{j}$ are shown in the top plot of Figure 7(b) (solid blue), together with the optimal threshold $\tau_{\rm opt}$ (horizontal dashed red). The $\rho({\bm{y}})$ error with $r=r^{*}=5$ as a function of $\tau$ can be seen in Figure 8(a) (right plot, solid blue), with its minimum indicated by the red square marker. This minimum is close to the minimum of the inaccessible $\rho({\bm{y}}_{\rm t})$ error (dashed gray). Here, as in the numerical examples of Section 5, we also observe the synchronous behavior of the feature removal errors on the training and test datasets.
@@
The removed ‘improper’ predictors are indicated with vertical solid light-gray lines in Figure 7(b) (middle and bottom). We have also extracted a few typical predictor data (colored vertical lines in Figure 7(a)) and displayed them in Figure 7(b) using the same colors and line styles for the corresponding polynomial fits. The retained predictors are in the left plot of Figure 7(b) and the removed predictors are in the right plot. It is easy to recognize the cone-type data, similar to the data in Figure 1 (right), in one of the discarded predictors (blue). All predictor data in the discarded gray zone around the dashed blue line in Figure 7(a) have this ‘conical’ shape. This means that the NIR data in this frequency band is situated on a hyper-cone. Even if this frequency band is not required to build a high-quality predictive IR model, it may become useful for testing other higher-dimensional manifold and nonlinear models.
@@
Finally, in Figure 8(b) we display the standard quality-of-prediction scatter plot which compares the exact values of $\bm{y}$} against their predictions $\hat{\bm{y}}$. As can be seen, the removal of the ‘improper’ predictors significantly improves the quality of prediction. In fact, the prediction error $\rho({\bm{y}}_{\rm t})$ on the test dataset goes down from $0.28$ to $0.05$ after the predictor removal procedure.
@@ 7 Conclusions
[…]
This work was funded by HZPC Research B.V., Averis Seeds B.V., BO-Akkerbouw, and European Agricultural Fund for Rural Development